# Microwave-driven multistability in a strongly interacting Rydberg atoms

### v1

@@
Microwave-driven multistability in a strongly interacting Rydberg atoms
Yu Ma1,2,⋆    Bang Liu1,2,⋆    Li-Hua Zhang1,2,⋆    Ya-Jun Wang1,2    Zheng-Yuan Zhang1,2    Shi-Yao Shao1,2    Qing Li1,2    Han-Chao Chen1,2    Jun Zhang1,2    Tian-Yu Han1,2    Qi-Feng Wang1,2    Jia-Dou Nan1,2    Yi-Ming Ying1,2    Dong-Yang Zhu1,2    Bao-Sen Shi1,2    Dong-Sheng Ding1,2,† 1Key Laboratory of Quantum Information, University of Science and Technology of China, Hefei, Anhui 230026, China. 2Synergetic Innovation Center of Quantum Information and Quantum Physics, University of Science and Technology of China, Hefei, Anhui 230026, China.
(August 20, 2024)
The interactions between Rydberg atoms and microwave fields provide a valuable framework for studying the complex dynamics out of equilibrium, exotic phases, and critical phenomena in many-body physics. This unique interplay allows us to explore various regimes of nonlinearity and phase transitions. Here, we observe a phase transition from the state in the regime of bistability to that in multistability in a strongly interacting Rydberg atoms by varying the microwave field intensity, accompanying with the breaking of $\mathbb{Z}_{3}$-symmetry. During the phase transition, the system experiences a hidden critical point, in which the multistable states are difficult to be identified. Through changing the initial state of system, we can identify a hidden multistable state and reveal a hidden trajectory of phase transition, allowing us to track to a hidden critical point. In addition, we observe multiple phase transitions in spectra, suggesting higher-order symmetry breaking. The reported results shed light on manipulating multistability in dissipative Rydberg atoms systems and hold promise in the applications of non-equilibrium many-body physics.

### v2

# Folded multistability and hidden critical point in microwave-driven Rydberg atoms

@@
Folded multistability and hidden critical point in microwave-driven Rydberg atoms
Yu Ma1,2,⋆    Bang Liu1,2,⋆    Li-Hua Zhang1,2,⋆    Ya-Jun Wang1,2    Zheng-Yuan Zhang1,2    Shi-Yao Shao1,2    Qing Li1,2    Han-Chao Chen1,2    Jun Zhang1,2    Tian-Yu Han1,2    Qi-Feng Wang1,2    Jia-Dou Nan1,2    Yi-Ming Yin1,2    Dong-Yang Zhu1,2    Bao-Sen Shi1,2    Dong-Sheng Ding1,2,† 1Key Laboratory of Quantum Information, University of Science and Technology of China, Hefei, Anhui 230026, China. 2Synergetic Innovation Center of Quantum Information and Quantum Physics, University of Science and Technology of China, Hefei, Anhui 230026, China.
(September 9, 2024)
The interactions between Rydberg atoms and microwave fields provide a valuable framework for studying the complex dynamics out of equilibrium, exotic phases, and critical phenomena in many-body physics. This unique interplay allows us to explore various regimes of nonlinearity and phase transitions. Here, we observe a phase transition from the state in the regime of bistability to that in multistability in strongly interacting Rydberg atoms by varying the microwave field intensity, accompanying with the breaking of $\mathbb{Z}_{3}$-symmetry. During the phase transition, the system experiences a hidden critical point, in which the multistable states are difficult to be identified. Through changing the initial state of system, we can identify a hidden multistable state and reveal a hidden trajectory of phase transition, allowing us to track to a hidden critical point. In addition, we observe multiple phase transitions in spectra, suggesting higher-order symmetry breaking. The reported results shed light on manipulating multistability in dissipative Rydberg atoms systems and hold promise in the applications of non-equilibrium many-body physics.